OpenStudy (anonymous):

What are the solutions of the equation: z2 - 12z + 36 = 0? A. -6, 6 B. 6, 6 C. 6, -6 D. -6, -6

OpenStudy (zzr0ck3r):

you can plug in the solutions and see if they work, or you can factor it

OpenStudy (zzr0ck3r):

what do you want to do?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

I am a parent trying to correct this test my Son turn in. I think he is cheating so I just want the right answer so I can bust him... He out of hand

OpenStudy (anonymous):

6,6

OpenStudy (anonymous):

thank you Passive

OpenStudy (anonymous):

how did you get that

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

ok

OpenStudy (anonymous):

I have 2 more question if you don't mind zz

OpenStudy (zzr0ck3r):

we are not allowed to just give answers....

OpenStudy (anonymous):

oh ok I will tell you what he put then you can tell me??

OpenStudy (anonymous):

If an object is dropped from a height of 144 feet, the function h(t)= -16t^2+144 gives the height of the object after t seconds. When will the object hit the ground? A. 9 s B. 6 s C. 1.5 s D. 3 s

OpenStudy (anonymous):

12z=z^2+36 Z^2 -12Z + 36 = 0 Z^2 -12Z + 36 = 0 ( Z - 6)^2 = 0 6,6